OpenStudy (anonymous):

I'm so lost on this question factor the trinomial 15b^2-22bc-5c^2?

OpenStudy (jdoe0001):

I'm thinking you're missing something there, recheck your polynomial

OpenStudy (anonymous):

yeah it was a term missing 15b^2-22bc-5c^2

OpenStudy (jdoe0001):

recheck it again :)

OpenStudy (anonymous):

factor the trinomial 15b^2-22bc-5c^2

OpenStudy (jdoe0001):

I see nothing there but just grouping, which isn't much of a factoring :|

OpenStudy (jdoe0001):

the negativity of the last term pretty much cancels it out as a perfect square, so :|

OpenStudy (anonymous):

15*-5=-75 25-3=22 25*-3=-75 \[15b ^{2}-22bc-5c ^{2}=15b ^{2}-\left( 25-3 \right)bc-5c ^{2}\] \[=15b ^{2}-25bc+3bc-5c ^{2}=5b \left( 3b-5c \right)+c \left( 3b-5c \right)=\left( 3b-5c \right)\left( 5b+c \right)\]